With a 3-fold increase in the radius of the circular orbit of an artificial Earth

satellite, the period of its rotation increases by 6 times. How many times will

the satellite's orbital speed change?


1
Expert's answer
2021-06-08T09:37:08-0400

as we know ,


"speed ={distance\over time}"


in circular orbit ,the relation between T,R,V is:


"V={2\pi R\over T}"


as given radius increased 3 times and time period by 6 times.

we get:


"V'={2\pi (3R)\over(6T)}={2 \pi R \over2T}"


"V'={V\over2}"


therefore speed will reduced to half the original.